# Dynamics of Cosmic Superstrings and the Overshoot Problem

### v1

@@ III.2 Case without radiation
In the case of D3-strings, $X^{2}\simeq\beta^{2}=1/9$ is not low enough to avoid overshooting, and the system has an evolution which is qualitatively similar to the F-string case, as can be seen in Fig. 3.
The case of NS5-strings as a loop fluid is instead quite promising for avoiding overshooting. In fact, as can be seen in Fig. 3, for a sufficiently high fraction of energy density in loops, the field does not overshoot.
Figure 3: Evolution of $\Phi$ from $\Phi_{0}=6M_{p}$ towards the minimum of the potential (9) located at $\Phi_{\rm min}=19M_{p}$ for the 3 values of $\beta$. For each $\beta$, 4 initial densities of loops have been tested: $Z_{0}^{2}=\{10^{-4},4\times 10^{-4},1.2\times 10^{-3},6.4\times 10^{-3}\}$. For $\beta=1/2$ and $\beta=1/3$, all of them overshoot, while for $\beta=1/6$ the trajectories with the higher values of $Z_{0}^{2}$ do not overshoot.
We can then study the evolution of the fractions of the energy densities $\Omega_{\rm k}$, $\Omega_{V}$, and $\Omega_{\rm loop}$ along the solutions of the autonomous system (24)-(25) with $\beta=1/6$ which do not overshoot. An illustrative result is shown in Fig. 4.
@@
Here, the initial conditions for the energy densities have been chosen keeping in mind a scenario where the volume modulus starts its evolution right at the end of inflation, so that, as it thaws, all its energy density is originally in its potential. This can happen in hybrid inflation models as brane-antibrane inflation Burgess et al. (2001); Dvali et al. (2001); Kachru et al. (2003a); Baumann et al. (2008, 2007); Cicoli et al. (2025), where string loops can be produced at the end of inflation when the volume acts as a waterfall field.
Note that, even for a small initial fraction of energy density in loops, the field does not overshoot. There are a few other interesting features of this case. First of all, during the evolution of the system, the fraction of energy density in loops grows significantly, reaching values that are even higher than the ones found in Sánchez González et al. (2025). The reason is the lower value of $\beta$: the system approaches the loop tracker fixed point $\mathcal{L}$ where $\Omega_{\rm loop}=1-\beta^{2}=35/36\simeq 0.972$, i.e. where the energy density in effective string loops can be as high as $97.2\%$!
What happens after the field has reached its minimum and starts oscillating is perhaps even more interesting. As shown in Fig. 4, the fraction of energy density in loops stabilises to about $50\%$! The reason for this is the following. As seen in (10), $\rho_{\rm loop}$ redshifts as matter plus the additional factor due to the decreasing tension. However, as soon as the field reaches its minimum, the tension starts oscillating around a fixed value, i.e. it stops decreasing. This means that from this point onwards $\rho_{\rm loop}\sim a^{-3}$, which is the same scaling behaviour of a field oscillating around its minimum, $\rho_{\phi}\sim a^{-3}$. This implies that the fraction of the energy densities remains on average constant:
$\frac{\Omega_{\rm loop}}{\Omega_{\rm k}+\Omega_{V}}\equiv\frac{Z^{2}}{X^{2}+Y^{2}}\simeq\text{const}\,.$ (35)
@@
III.3 Multiple string loop species
@@
$\rho_{F}\propto a^{-3}\mathcal{V}^{-1/2}=a^{-3}e^{-\frac{1}{2}\sqrt{\frac{3}{2}}\Phi}\,$, (36)
@@
$\rho_{3}\propto a^{-3}\mathcal{V}^{-1/3}=a^{-3}e^{-\frac{1}{3}\sqrt{\frac{3}{2}}\Phi}\,$, (37)
@@
$\rho_{5}\propto a^{-3}\mathcal{V}^{-1/6}=a^{-3}e^{-\frac{1}{6}\sqrt{\frac{3}{2}}\Phi}.$ (38)
@@
III.4 Case with radiation
If we include a radiation background fluid with equation of state $p_{\rm f}=\frac{1}{3}\rho_{\rm f}$, the dynamics of the system becomes even richer. Recalling that $\Omega_{\rm f}\equiv W^{2}$ and including the continuity equation of the fluid in the dynamical system, we get:
$X^{\prime}$ $=\frac{3}{2}\bigg[X\left(X^{2}-Y^{2}+\frac{W^{2}}{3}-1\right)$
$\hskip 26.74565pt+\beta\left(1-X^{2}-W^{2}\right)+(\lambda-\beta)Y^{2}\bigg],$ (39)
$Y^{\prime}$ $=\frac{3}{2}Y\left(X^{2}-Y^{2}+\frac{W^{2}}{3}-\lambda X+1\right),$ (40)
$W^{\prime}$ $=\frac{3}{2}W\left[X^{2}-Y^{2}+\frac{1}{3}\left(W^{2}-1\right)\right].$ (41)
@@ III.4 Case with radiation
$\mathcal{S}$ $\frac{4}{3\lambda}$ $\frac{2\sqrt{2}}{3\lambda}$ 0 $\sqrt{1-\frac{8}{3\lambda^{2}}}$ $\qquad\qquad\qquad\quad\lambda\geq 2\sqrt{\frac{2}{3}}\,$, $\beta\neq\frac{\lambda}{4}$
$\mathcal{LR}$ $\frac{4}{3\lambda}$ $y_{\rm lr}$ $\sqrt{\frac{32}{9\lambda^{2}}-4y^{2}_{\rm lr}}$ $\sqrt{1-\frac{16}{3\lambda^{2}}+3y_{\rm lr}^{2}}$ $\qquad\qquad\qquad\quad\lambda\geq 2\sqrt{\frac{2}{3}}\,$, $\beta=\frac{\lambda}{4}$
$\mathcal{T}_{2}$ $\frac{1}{3\beta}$ 0 $\frac{\sqrt{2}}{3\beta}$ $\sqrt{1-\frac{1}{3\beta^{2}}}$ $\qquad\qquad\qquad\qquad\;\forall\lambda,\,\beta\geq\frac{1}{\sqrt{3}}$
$\mathcal{F}$ 0 0 0 1 $\qquad\qquad\qquad\qquad\,\forall\lambda$ and $\forall\beta$
Table 2: Fixed points (FP) of the autonomous system (39)-(40)-(41) with existence conditions. These all exist together with the fixed points of Tab. 1.
Once again, the fixed points for this system in the case of an exponential potential, i.e. $\lambda=\text{const}$, have been catalogued in Sánchez González et al. (2025) and we report them in Tab. 2.
All the fixed points of the dynamical system in the absence of radiation (24)-(25) are still fixed points of this system, as they belong to the $W=0$ section of the phase space, albeit with altered stability properties. The new fixed points of Tab. 2 are exclusive to the case with radiation. In fact, we have the fluid dominated fixed point $\mathcal{F}$ in which all of the energy density of the universe is in radiation. In the scaling solution $\mathcal{S}$, or radiation tracker, also analysed in Copeland et al. (1998); Brunelli et al. (2025), the loops do not play any role. Then we have two new tracker solutions, namely the loop-radiation tracker $\mathcal{LR}$ and a second mixed tracker $\mathcal{T}_{2}$.
The fixed point $\mathcal{LR}$ exists only in the specific case where $\beta=\lambda/4$, and is parametrised by its $Y$ coordinate, $y_{\rm lr}$. In fact, this is not a single fixed point but a line in phase space with the $Y$ coordinate lying in the range $y_{\rm min}\leq y_{\rm lr}\leq\frac{2\sqrt{2}}{3\lambda}$, with $y_{\rm min}$ depending on $\lambda$ as:
$y_{\rm min}=\begin{cases}\sqrt{\frac{16}{9\lambda^{2}}-\frac{1}{3}}\quad&\text{for }\,2\sqrt{\frac{2}{3}}\leq\lambda\leq\frac{4}{\sqrt{3}}\\ 0\quad&\text{for }\,\lambda\geq\frac{4}{\sqrt{3}}\end{cases}\,.$ (42)
We note that this fixed point plays no role in the following discussion since we lack a concrete realisation of effective strings with $\beta=3/4$ (for $\lambda=3$).
@@
Figure 6: Stability map of the fixed points of the dynamical system (39)-(41) for an exponential potential.
The exponential approximation (26) of the LVS potential in the runaway region, in the case of F-strings, corresponds to the point $(\beta=1/2,\lambda=3)$ in parameter space, yielding the loop tracker $\mathcal{L}$ as the final attractor, in agreement with the results of Sánchez González et al. (2025). However, if the field reaches the minimum before the system approaches $\mathcal{L}$, there is a chance of avoiding overshooting. In fact, for a purely exponential potential without a minimum, the system will first hover around the radiation tracker fixed point $\mathcal{S}$ and then evolve towards the final attractor $\mathcal{L}$, as can be appreciated in Fig. 7.
@@
As shown in Fig. 7, the system evolves first towards a kination epoch, then lingers around the radiation tracker for a few efolds and finally reaches the minimum as it is evolving towards $\mathcal{L}$. Thus $\Omega_{\rm loop}=Z^{2}$ starts growing before the field starts oscillating, but without overcoming the dominant radiation. An interesting feature of this system emerges as $\Phi$ starts oscillating around its minimum. At this stage, the cosmic string loops behave (on average) as matter and satisfy (35). However, now the system includes also radiation which redshifts faster than matter, as $\rho_{\rm rad}\sim a^{-4}$. Therefore, the ratios $Z^{2}/W^{2}$ and $(X^{2}+Y^{2})/W^{2}$ both grow during the oscillations. This implies that, as $W^{2}$ decreases, $X^{2}+Y^{2}$ increases, so does $Z^{2}$ in order for (35) to be always satisfied. The resulting effect is a growth in the fraction of the energy density in loops as the the volume modulus oscillates around minimum. Just like in the case with no radiation, $\Omega_{\rm loop}=Z^{2}$ can reach significant values, around $50\%$, as can be seen from Fig. 8. As radiation redshifts away, this fraction settles and we are back to the case without radiation.
A similar behaviour is expected for effective strings. However, as argued in Sec. III.3, for $\beta<1/2$ the energy density of loops redshifts slower with respect to the F-string case, and so the loop tracker is reached faster. Thus, for the initial fraction of energy densities used to plot Fig. 8, D3-strings with $\beta=1/3$ will lead the field to overshoot. However, if the hierarchy between $W_{0}^{2}$ and $Z_{0}^{2}$ is higher, the system will mimic precisely what happens in Fig. 8. The evolution of the energy densities in a case with $\beta=1/3$ where the field does not overshoot is shown in Fig. 9.
@@
III.5 Dissipation due to GW emission
The present analysis has neglected the loss in energy density of loops due to gravitational wave emission, which is a continuous process. An argument for which one should be careful in doing this is the following, adapted from Sánchez González et al. (2025). The typical power of emission of gravitational wave radiation is given by:
$P_{\scriptscriptstyle GW}\simeq\frac{\Gamma}{8\pi}\left(\frac{\mu}{M_{p}}\right)^{2}\,$, (43)
@@ III.5 Dissipation due to GW emission
$\tau_{\scriptscriptstyle GW}\simeq 8\pi\frac{\ell M_{p}^{2}}{\Gamma\mu}\,.$ (44)
While for F-strings, depending on initial conditions, the argument of Sánchez González et al. (2025) shows that the loss is essentially negligible in a Hubble time, for effective strings things are different. In particular, for NS5-strings, the tension goes as $\mu\sim M_{s}^{2}\mathcal{V}^{2/3}\sim M_{p}^{2}/\mathcal{V}^{1/3}$, so that:
$\tau_{\scriptscriptstyle GW}\sim 8\pi\frac{\mathcal{V}^{1/3}}{\Gamma}\,\ell\,.$ (45)
@@
$\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}\simeq 8\pi\frac{\sqrt{\mathcal{V}}}{\Gamma}\,\frac{H}{M_{p}}\,.$ (46)
@@
Depending on the initial values of $H$ and $\mathcal{V}$, this fraction may or may not be relevant as the field starts evolving. What is surely interesting is that, unlike the case of F-strings, the emission process becomes more and more relevant as the volume approaches its minimum. A simple way to see this, is tracking the behaviour of the ratio (46) along the evolution of the system. As the volume rolls down its potential, it quickly enters a kinating epoch. During such a period, $\mathcal{V}\sim a^{3}\sim e^{3N}$, while, on the other hand, $H\sim a^{-3}\sim e^{-3N}$. Therefore, if the kinating epoch lasts $N_{\rm k}$ efolds:
$\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}=\left(\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}\right)_{\rm in}e^{-\frac{3}{2}N_{\rm k}}\,.$ (47)
@@
After that, the system enters the loop tracking regime. One can use the second Friedman equation (16) with $\rho_{\rm rad}=0$ to retrieve how $\mathcal{V}$ and $H$ evolve with the scale factor in that period. For NS5-strings ($\beta=1/6$), one gets $\mathcal{V}\sim a^{1/2}\sim e^{N/2}$, while $H\sim a^{-\frac{37}{24}}\sim e^{-\frac{37}{24}N}$. Supposing the transition between kination and loop tracking is instantaneous, and the loop-tracker period lasts $N_{\rm loop}$ efolds, one finds:
$\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}=\left(\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}\right)_{\rm in}e^{-\frac{3}{2}N_{\rm k}-\frac{31}{24}N_{\rm loop}}\,$, (48)
@@
implying that the loss of energy into GWs becomes increasingly relevant at late times. This is to be contrasted with Sánchez González et al. (2025) where the focus on F-strings implied that GW production was more efficient at early times.
@@
By neglecting energy loss into GWs, we are assuming that the parameters in the above equation, in particular the initial string length and compactification volume, are such that $\tau_{\scriptscriptstyle GW}\gg t_{H}$ at all times. If that is not the case, the dependence of the loop size on the GW emission has to be included. As argued in Ghoshal et al. (2025), the size of the loop decreases due to GW emission. This effect may be more or less efficient than the growth of the string length due to the decrease in tension. For NS5-strings, the effect can be much larger at late times, implying that the physical string length could decrease in time, further reducing $\tau_{\scriptscriptstyle GW}$. An efficient emission of GWs would actually help in solving the overshoot problem, as the emitted radiation contributes in slowing the field down as it rolls towards its minimum. On the other hand, the fraction of residual energy density in loops at the minimum would be surely affected. Therefore, to assess the late-time evolution of the system, a more detailed analysis including the energy loss should be carried out. Such a study should take into account the dependence of the size of the loops on the produced radiation, which influences both the frequency of the emitted GWs and their amplitude. A detailed analysis of the GW spectrum for a network of cosmic superstrings has recently been conducted in Ghoshal et al. (2025), in a similar stringy embedding as the present paper. We leave the precise analysis of the efficiency of the production and the GW spectrum of the setup presented here for future work.
IV Oscillating tension

### v2

@@ 3.2 Case without radiation
In the case of D3-strings, $X^{2}\simeq\beta^{2}=1/9$ is not low enough to avoid overshooting, and the system has an evolution which is qualitatively similar to the F-string case, as can be seen in Fig. 3. The case of NS5-strings as a loop fluid is instead quite promising for avoiding overshooting. In fact, as can be seen in Fig. 3, for a sufficiently high fraction of energy density in loops, the field does not overshoot.
Figure 3: Evolution of $\Phi$ from $\Phi_{0}=6M_{p}$ towards the minimum of the potential (3.2) located at $\Phi_{\rm min}=19M_{p}$ for the 3 values of $\beta$. For each $\beta$, 4 initial densities of loops have been tested: $Z_{0}^{2}=\{10^{-4},4\times 10^{-4},1.2\times 10^{-3},6.4\times 10^{-3}\}$. For $\beta=1/2$ and $\beta=1/3$, all of them overshoot, while for $\beta=1/6$ the trajectories with the higher values of $Z_{0}^{2}$ do not overshoot.
In Fig. 4 we perform a systematic scan over the $(\beta,\Omega_{\rm loop}^{0})$ parameter space, in order to determine the values of these parameters that do not lead to volume modulus overshoot. In this analysis all other parameters are kept fixed, in particular $\Phi_{0}=6M_{p}$, $\dot{\Phi}_{0}=0$ and $\Phi_{\rm min}=19M_{p}$. We observe that a non-vanishing energy density in string loops is always required to avoid overshooting and that this fraction increases with increasing $\beta$. Furthermore we observe that overshoot is inevitable for $\beta>0.23$, in accordance with the findings in Fig. 3.
Figure 4: Systematic scan over the $(\beta,\Omega_{\rm loop}^{0})$ space. The blue region corresponds to values of the parameters for which the field does not overshoot its minimum while red region represents the parameters for which it does. The other initial conditions are fixed as: $\Phi_{0}=6$, $\dot{\Phi}_{0}=0$ and $\Phi_{\rm min}=19$ Left: Scan over all the meaningful values of $\beta$, up to $\beta=1/2$. The plot shows that the field will always overshoot for $\beta>0.23$. Right: Zoom over the $\beta<0.2$ region, showing how the minimal $\Omega_{\rm loop}^{0}$ grows exponentially with $\beta$.
We can then study the evolution of the fractions of the energy densities $\Omega_{\rm k}$, $\Omega_{V}$, and $\Omega_{\rm loop}$ along the solutions of the autonomous system (3.17)-(3.18) with $\beta=1/6$ which do not overshoot. An illustrative result is shown in Fig. 5.
@@
Here, the initial conditions for the energy densities have been chosen keeping in mind a scenario where the volume modulus starts its evolution right at the end of inflation, so that, as it thaws, all its energy density is originally in its potential. This can happen in hybrid inflation models as brane-antibrane inflation [13, 25, 29, 7, 6, 16], where string loops can be produced at the end of inflation when the volume acts as a waterfall field.
Note that, even for a small initial fraction of energy density in loops, the field does not overshoot. There are a few other interesting features of this case. First of all, during the evolution of the system, the fraction of energy density in loops grows significantly, reaching values that are even higher than the ones found in [36]. The reason is the lower value of $\beta$: the system approaches the loop tracker fixed point $\mathcal{L}$ where $\Omega_{\rm loop}=1-\beta^{2}=35/36\simeq 0.972$, i.e. where the energy density in effective string loops can be as high as $97.2\%$!
What happens after the field has reached its minimum and starts oscillating is perhaps even more interesting. As shown in Fig. 5, the fraction of energy density in loops stabilises to about $50\%$! The reason for this is the following. As seen in (3.3), $\rho_{\rm loop}$ redshifts as matter plus the additional factor due to the decreasing tension. However, as soon as the field reaches its minimum, the tension starts oscillating around a fixed value, i.e. it stops decreasing. This means that from this point onwards $\rho_{\rm loop}\sim a^{-3}$, which is the same scaling behaviour of a field oscillating around its minimum, $\rho_{\phi}\sim a^{-3}$. This implies that the fraction of the energy densities remains on average constant:
$\frac{\Omega_{\rm loop}}{\Omega_{\rm k}+\Omega_{V}}\equiv\frac{Z^{2}}{X^{2}+Y^{2}}\simeq\text{const}\,.$ (3.28)
@@
3.3 Multiple string loop species
@@
$\rho_{F}\propto a^{-3}\mathcal{V}^{-1/2}=a^{-3}e^{-\frac{1}{2}\sqrt{\frac{3}{2}}\Phi}\,$, (3.29)
@@
$\rho_{3}\propto a^{-3}\mathcal{V}^{-1/3}=a^{-3}e^{-\frac{1}{3}\sqrt{\frac{3}{2}}\Phi}\,$, (3.30)
@@
$\rho_{5}\propto a^{-3}\mathcal{V}^{-1/6}=a^{-3}e^{-\frac{1}{6}\sqrt{\frac{3}{2}}\Phi}.$ (3.31)
@@
3.4 Case with radiation
If we include background fluid with equation of state $p_{\rm f}=\omega\,\rho_{\rm f}$, the dynamics of the system becomes even richer. Recalling that $\Omega_{\rm f}\equiv W^{2}$ and including the continuity equation of the fluid in the dynamical system, we get:
$X^{\prime}$ $=\frac{3}{2}\bigg[X\left(X^{2}-Y^{2}+\omega W^{2}-1\right)+\beta\left(1-X^{2}-W^{2}\right)+(\lambda-\beta)Y^{2}\bigg],$ (3.32)
$Y^{\prime}$ $=\frac{3}{2}Y\left(X^{2}-Y^{2}+\omega W^{2}-\lambda X+1\right),$ (3.33)
$W^{\prime}$ $=\frac{3}{2}W\left[X^{2}-Y^{2}+\omega\left(W^{2}-1\right)\right].$ (3.34)
@@ 3.4 Case with radiation
$\mathcal{S}$ $\frac{\omega+1}{\lambda}$ $\frac{\sqrt{1-\omega^{2}}}{\lambda}$ 0 $\sqrt{1-\frac{2(\omega+1)}{\lambda^{2}}}$ $\lambda\geq\sqrt{2(\omega+1)}\,$, $\beta\neq\lambda\frac{\omega}{(\omega+1)}$
$\mathcal{LR}$ $\frac{\omega+1}{\lambda}$ $y_{\rm lr}$ $\sqrt{\frac{(1-\omega^{2})(1+\omega)}{\omega\lambda^{2}}-\frac{\omega+1}{\omega}y^{2}_{\rm lr}}$ $\sqrt{1-\frac{(\omega+1)^{2}}{\omega\lambda^{2}}+\frac{y_{\rm lr}^{2}}{\omega}}$ $\lambda\geq\sqrt{2(\omega+1)}$, $\beta=\lambda\frac{\omega}{(\omega+1)},\,\omega>0$
$\mathcal{T}_{2}$ $\frac{\omega}{\beta}$ 0 $\frac{\sqrt{\omega(1-\omega)}}{\beta}$ $\sqrt{1-\frac{\omega}{\beta^{2}}}$ $\qquad\qquad\forall\lambda,\,\beta\geq\sqrt{\omega}$
$\mathcal{F}$ 0 0 0 1 $\qquad\qquad\,\forall\lambda$ and $\forall\beta$
Table 2: Fixed points (FP) of the autonomous system (3.32)-(3.33)-(3.34) with existence conditions. These all exist together with the fixed points of Tab. 1.
Once again, the fixed points for this system in the case of an exponential potential, i.e. $\lambda=\text{const}$ in the case of a radiation fluid $\omega=1/3$, have been catalogued in [36], and we report them in Tab. 2. All the fixed points of the dynamical system in the absence of fluid (3.17)-(3.18) are still fixed points of this system, as they belong to the $W=0$ section of the phase space, albeit with altered stability properties. The new fixed points of Tab. 2 are exclusive to the case with radiation. In fact, we have the fluid dominated fixed point $\mathcal{F}$ in which all of the energy density of the universe is in radiation. In the scaling solution $\mathcal{S}$, or tracker, also analysed in [24, 10], the loops do not play any role. Then we have two new tracker solutions, namely the loop-radiation tracker $\mathcal{LR}$ and a second mixed tracker $\mathcal{T}_{2}$.
The fixed point $\mathcal{LR}$ exists only in the specific case where $\beta=\lambda\omega/(\omega+1)$, and is parametrised by its $Y$ coordinate, $y_{\rm lr}$. In fact, this is not a single fixed point but a line in phase space with the $Y$ coordinate lying in the range $y_{\rm min}\leq y_{\rm lr}\leq\frac{\sqrt{1-\omega^{2}}}{\lambda}$, with $y_{\rm min}$ depending on $\lambda$ as:
$y_{\rm min}=\begin{cases}\sqrt{\frac{(\omega+1)^{2}}{\lambda^{2}}-\omega}\quad&\text{for }\,\sqrt{2(\omega+1)}\leq\lambda\leq\frac{\omega+1}{\sqrt{\omega}}\\ 0\quad&\text{for }\,\lambda\geq\frac{\omega+1}{\sqrt{\omega}}\end{cases}\,.$ (3.35)
We note that this fixed point plays no role in the following discussion since we lack a concrete realisation of effective strings with $\beta=3/4$ (for $\lambda=3$, $\omega=1/3$).
@@
From this point onward, we focus on the case where the background fluid is radiation, $\omega=1/3$, as this is the most relevant case.
@@
Figure 7: Stability map of the fixed points of the dynamical system (3.32)-(3.34) for an exponential potential and a radiation background fluid.
The exponential approximation (3.19) of the LVS potential in the runaway region, in the case of F-strings, corresponds to the point $(\beta=1/2,\lambda=3)$ in parameter space, yielding the loop tracker $\mathcal{L}$ as the final attractor, in agreement with the results of [36]. However, if the field reaches the minimum before the system approaches $\mathcal{L}$, there is a chance of avoiding overshooting. In fact, for a purely exponential potential without a minimum, the system will first hover around the radiation tracker fixed point $\mathcal{S}$ and then evolve towards the final attractor $\mathcal{L}$, as can be appreciated in Fig. 8.
@@
As shown in Fig. 8, the system evolves first towards a kination epoch, then lingers around the radiation tracker for a few efolds and finally reaches the minimum as it is evolving towards $\mathcal{L}$. Thus $\Omega_{\rm loop}=Z^{2}$ starts growing before the field starts oscillating, but without overcoming the dominant radiation. An interesting feature of this system emerges as $\Phi$ starts oscillating around its minimum. At this stage, the cosmic string loops behave (on average) as matter and satisfy (3.28). However, now the system includes also radiation which redshifts faster than matter, as $\rho_{\rm rad}\sim a^{-4}$. Therefore, the ratios $Z^{2}/W^{2}$ and $(X^{2}+Y^{2})/W^{2}$ both grow during the oscillations. This implies that, as $W^{2}$ decreases, $X^{2}+Y^{2}$ increases, so does $Z^{2}$ in order for (3.28) to be always satisfied. The resulting effect is a growth in the fraction of the energy density in loops as the the volume modulus oscillates around minimum. Just like in the case with no radiation, $\Omega_{\rm loop}=Z^{2}$ can reach significant values, around $50\%$, as can be seen from Fig. 9. As radiation redshifts away, this fraction settles and we are back to the case without radiation.
A similar behaviour is expected for effective strings. However, as argued in Sec. 3.3, for $\beta<1/2$ the energy density of loops redshifts slower with respect to the F-string case, and so the loop tracker is reached faster. Thus, for the initial fraction of energy densities used to plot Fig. 9, D3-strings with $\beta=1/3$ will lead the field to overshoot. However, if the hierarchy between $W_{0}^{2}$ and $Z_{0}^{2}$ is higher, the system will mimic precisely what happens in Fig. 9. The evolution of the energy densities in a case with $\beta=1/3$ where the field does not overshoot is shown in Fig. 10.
@@
3.5 Dissipation due to GW emission
The present analysis has neglected the loss in energy density of loops due to gravitational wave emission, which is a continuous process. An argument for which one should be careful in doing this is the following, adapted from [36]. The typical power of emission of gravitational wave radiation is given by:
$P_{\scriptscriptstyle GW}\simeq\frac{\Gamma}{8\pi}\left(\frac{\mu}{M_{p}}\right)^{2}\,$, (3.36)
@@ 3.5 Dissipation due to GW emission
$\tau_{\scriptscriptstyle GW}\simeq 8\pi\frac{\ell M_{p}^{2}}{\Gamma\mu}\,.$ (3.37)
While for F-strings, depending on initial conditions, the argument of [36] shows that the loss is essentially negligible in a Hubble time, for effective strings things are different. In particular, for NS5-strings, the tension goes as $\mu\sim M_{s}^{2}\mathcal{V}^{2/3}\sim M_{p}^{2}/\mathcal{V}^{1/3}$, so that:
$\tau_{\scriptscriptstyle GW}\sim 8\pi\frac{\mathcal{V}^{1/3}}{\Gamma}\,\ell\,.$ (3.38)
@@
$\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}\simeq 8\pi\frac{\sqrt{\mathcal{V}}}{\Gamma}\,\frac{H}{M_{p}}\,.$ (3.39)
@@
Depending on the initial values of $H$ and $\mathcal{V}$, this fraction may or may not be relevant as the field starts evolving. What is surely interesting is that, unlike the case of F-strings, the emission process becomes more and more relevant as the volume approaches its minimum. A simple way to see this, is tracking the behaviour of the ratio (3.39) along the evolution of the system. As the volume rolls down its potential, it quickly enters a kinating epoch. During such a period, $\mathcal{V}\sim a^{3}\sim e^{3N}$, while, on the other hand, $H\sim a^{-3}\sim e^{-3N}$. Therefore, if the kinating epoch lasts $N_{\rm k}$ efolds:
$\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}=\left(\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}\right)_{\rm in}e^{-\frac{3}{2}N_{\rm k}}\,.$ (3.40)
@@
After that, the system enters the loop tracking regime. One can use the second Friedman equation (3.9) with $\rho_{\rm rad}=0$ to retrieve how $\mathcal{V}$ and $H$ evolve with the scale factor in that period. For NS5-strings ($\beta=1/6$), one gets $\mathcal{V}\sim a^{1/2}\sim e^{N/2}$, while $H\sim a^{-\frac{37}{24}}\sim e^{-\frac{37}{24}N}$. Supposing the transition between kination and loop tracking is instantaneous, and the loop-tracker period lasts $N_{\rm loop}$ efolds, one finds:
$\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}=\left(\frac{\tau_{\scriptscriptstyle GW}}{t_{H}}\right)_{\rm in}e^{-\frac{3}{2}N_{\rm k}-\frac{31}{24}N_{\rm loop}}\,$, (3.41)
@@
implying that the loss of energy into GWs becomes increasingly relevant at late times. This is to be contrasted with [36] where the focus on F-strings implied that GW production was more efficient at early times.
@@
By neglecting energy loss into GWs, we are assuming that the parameters in the above equation, in particular the initial string length and compactification volume, are such that $\tau_{\scriptscriptstyle GW}\gg t_{H}$ at all times. If that is not the case, the dependence of the loop size on the GW emission has to be included. As argued in [27], the size of the loop decreases due to GW emission. This effect may be more or less efficient than the growth of the string length due to the decrease in tension. For NS5-strings, the effect can be much larger at late times, implying that the physical string length could decrease in time, further reducing $\tau_{\scriptscriptstyle GW}$. An efficient emission of GWs would actually help in solving the overshoot problem, as the emitted radiation contributes in slowing the field down as it rolls towards its minimum. On the other hand, the fraction of residual energy density in loops at the minimum would be surely affected. Therefore, to assess the late-time evolution of the system, a more detailed analysis including the energy loss should be carried out. Such a study should take into account the dependence of the size of the loops on the produced radiation, which influences both the frequency of the emitted GWs and their amplitude. A detailed analysis of the GW spectrum for a network of cosmic superstrings has recently been conducted in [27], in a similar stringy embedding as the present paper. We leave the precise analysis of the efficiency of the production and the GW spectrum of the setup presented here for future work.
4 Oscillating tension